OpenStudy (theamazing.phan):

if the radius, r, is doubled, how does it affect the area of a circle?

OpenStudy (skyvoltage):

MARKIPLIER!!!!!!!!!

Vocaloid (vocaloid):

hint: A = pi * r^2 if we double r we get A = pi * (2r)^2 = pi * 4 * r^2 so what changes?